OpenStudy (anonymous):

Convert the complex number 2 - 2root3 into its polar representation. ??

OpenStudy (anonymous):

@SolomonZelman @perl @Kainui

OpenStudy (anonymous):

@OOOPS @Oooky

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

\((2, -2\sqrt3)\), right?

OpenStudy (anonymous):

first off, x =2, y = -2root3, hence x^2+y^2 =r^2 =?

OpenStudy (anonymous):

I guide you how to find it out, I don't do your homework

OpenStudy (anonymous):

answer me step by step, you can get it, it is not hard x^2 +y^2 =?

OpenStudy (anonymous):

http://www.wolframalpha.com/widgets/view.jsp?id=47e6fd9e600778e368177b1edd70b025 try using this calculator

OpenStudy (anonymous):

right ok so x^2 is 4 and y^2 = 12?

OpenStudy (anonymous):

@Oooky on test, we don't have math tools

OpenStudy (anonymous):

so r = 4?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

It was for future reference :P

OpenStudy (anonymous):

now, take r out by factor (2, -2root3) what do you get?

OpenStudy (anonymous):

ok, may be this is hard for you, I guide you this :)

OpenStudy (anonymous):

yeah how do i do that haha

OpenStudy (anonymous):

good luck :)))

OpenStudy (anonymous):

for x =2 , do you agree that 2 = 4*\(\dfrac{2}{4}\)

OpenStudy (anonymous):

hence if I take 4 out, then the rest is 1/2, right?

OpenStudy (anonymous):

yup i got that

OpenStudy (anonymous):

do the same with y, we have y = \(-2\sqrt 3= 4*-2\sqrt3/4 \) if I take 4 out, I have the rest is \(-\dfrac{\sqrt 3}{2}\), right?

OpenStudy (anonymous):

so far, I have \(4(\dfrac{1}{2}, -\dfrac{\sqrt3}{2}) \) got me so far? I just take 4 out from the original (x, y)

OpenStudy (anonymous):

yeah I follow that

OpenStudy (anonymous):

cos ??= 1/2?

OpenStudy (anonymous):

pi/3?

OpenStudy (anonymous):

oh, there is an easier way to do, hihihi... but by mistake, I go on this way. go ahead or go other way?? hahahah.... but the step to get r is perfect,

OpenStudy (anonymous):

either way haha

OpenStudy (anonymous):

ok, \(tan \theta =\dfrac{y}{x}\), just put the original x, y in to get tan

OpenStudy (anonymous):

I get tan theta = root 3

OpenStudy (anonymous):

-root 3, right? not root 3

OpenStudy (anonymous):

that's pi/3

OpenStudy (anonymous):

oh right so -pi/3

OpenStudy (anonymous):

positive one, please

OpenStudy (anonymous):

so is it A?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sqrt[4]{3} \cos(60) = i \sin(60)\]

OpenStudy (anonymous):